OpenStudy (anonymous):

Establish the following identity: sin^2(-theta)+cos^2(-theta)=1 i understand sin^2(theta)+cos^2(theta)=1 and cos(-theta)=cos(theta) and sin(-theta)=-sin(theta) so how do i approach the identity above.

OpenStudy (binary3i):

see you have sin(-theta) = - sin(theta) and cos(-theta) = cos (theta) just substitute it into cos^2(-theta) + sin^2(-theta)=1 it will be cos^2(theta)+(-1)^2sin^2(theat)=1 cos^2(theta) + sin^2(theta)=1

OpenStudy (shamim):

no

OpenStudy (binary3i):

what no

OpenStudy (shamim):

\[\sin ^{2}(- \theta)=(-\sin \theta)^{2}=\sin ^{2} \theta\]

OpenStudy (shamim):

did u get it

OpenStudy (binary3i):

did you saw what i have written above. see properly.

OpenStudy (anonymous):

ohh okay what shamim put about sin^2(-theta) =sin^2(theta) just made my life a whole lot better thx lol. and u too binary